OpenStudy (anonymous):

MEDAL f(x)=(x^2+3x-4) and g(x)=(x+4) a) find FOG and state the domain b) find f/g and state the domain.

OpenStudy (anonymous):

@campbell_st @ganeshie8 @dan815 @Nnesha @wolf1728

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

HI! Please help me!

OpenStudy (misty1212):

\[f\circ g(x)=f(g(x))\] we work from the inside out

OpenStudy (misty1212):

\[f(g(x))=f(x+4)\] is a start

OpenStudy (anonymous):

Is for the first one x^2+11x+32 with the domain as all real numbers?

OpenStudy (misty1212):

then since \[f(\diamondsuit)=\diamondsuit ^2+3\diamondsuit-4\] you have \[f(x+4)=(x+4)^2+3(x+4)-4\]

OpenStudy (misty1212):

then algebra give me a sec but the domain is all real numbers for sure

OpenStudy (anonymous):

I mean +28 at the end

OpenStudy (misty1212):

i get \((x+8) (x+3)\) or \[x^2+11x+24\] same thing

OpenStudy (anonymous):

Okay, and fr the second one?

OpenStudy (misty1212):

oh hold on let me see about the constant again

OpenStudy (misty1212):

24 is right

OpenStudy (anonymous):

Okay and how do I do f/g

OpenStudy (misty1212):

at the risk of sounding silly, put f over g

OpenStudy (misty1212):

ok i copied that wrong it is \[\frac{x^2+3x-4}{x+4}\] so the domain is all numbers except \(-4\)

OpenStudy (misty1212):

now we can factor and cancel \[\frac{(x+4)(x-1)}{x+4}=x-1\]

OpenStudy (anonymous):

Now what?

OpenStudy (anonymous):

(x-1)/1 = (x-1)?

OpenStudy (misty1212):

now go have a nice glass of wine since we are done

OpenStudy (anonymous):

So how would I write the solution?

OpenStudy (anonymous):

y=x-1?

OpenStudy (misty1212):

as \(x-1\) for \(x\neq -4\)

OpenStudy (anonymous):

Thank you so much!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]